suppose that tge demand curve facing OPEC is given by P=120-2Q and that ecah members cpst of producing oil AC=MC =$20.Find the cartels profit maximizing total output and price

1
Expert's answer
2021-11-22T15:09:36-0500
"\pi=TR-TC"

"\pi=(120-2Q)Q-20Q"

"\pi=100Q-2Q^2"

"\frac{\delta \pi}{\delta Q}=100-4Q"

"Q=25"

"p=70"

"\pi=1250"